solve system of equation by elimination
2x+y+4z=1
3x+2y+6z=2
5x-3y-10z=5

Answer by nyc_function(2741)

Show some effort or else your questions will be ignored.
ANSWERS: x = 4/5, y = 1 and z = -2/5